### v1

# backgroundcolor=gray!3, roundcorner=3pt

Abstract
Human guidance is often desired in reinforcement learning to improve the performance of the learning agent. However, human insights are often mere opinions and educated guesses rather than well-formulated arguments. While opinions are subject to uncertainty, e.g., due to partial informedness or ignorance about a problem, they also emerge earlier than hard evidence could be produced. Thus, guiding reinforcement learning agents through opinions offers the potential for more performant learning processes, but comes with the challenge of modeling and managing opinions in a formal way. In this article, we present a method to guide reinforcement learning agents through opinions. To this end, we provide an end-to-end method to model and manage advisors’ opinions. To assess the utility of the approach, we evaluate it with synthetic and human advisors, at different levels of uncertainty, and under multiple advise strategies. Our results indicate that opinions, even if uncertain, improve the performance of reinforcement learning agents, resulting in higher rewards, more efficient exploration, and a better reinforced policy. Although we demonstrate our approach in a simplified topological running example, our approach is applicable to complex problems with higher dimensions as well.
keywords:
artificial intelligence, belief, domain-specific languages, guided reinforcement learning, human guidance, machine learning, opinion, subjective logic, uncertainty
1 Introduction
Reinforcement learning (RL) [48] is a machine learning paradigm in which an autonomous agent explores its surroundings and learns optimal behavior through trial and error. Because of the continuous learning process, RL is particularly well-suited to deal with open-ended problems that feature unforeseen situations, tasks, and environments. Thanks to these benefits, the popularity of RL has been steadily increasing in domains where dealing with complexity is essential, such as fine-tuning the control of manufacturing processes at runtime [14], inferring simulation components of digital twins [20], and generating repair actions for conceptual domain models [6].
However, autonomous behavior does not obviate the need for human agency [7]. Human input to autonomous systems often augments machine intelligence with higher-level strategic directives that require creative problem-solving skills [41]. RL is no exception to this rule either. The body of knowledge on guided RL methods is substantial and rapidly growing.
General guidance-based techniques allow humans to inform agents about future aspects of the task, e.g., interesting regions to explore [47] or a trajectory (i.e., a sequence of actions) to follow [50]. A severe limitation of such techniques is the inconsistent correctness of human advice. As Scherf et al [45] report, the majority of interactive or guided reinforcement learning approaches assume the human input to be useful and correct [36]. However, this is not always true in real applications [31, 34, 4]. Informed advisors, such as experts, often express advice in the form of opinions—cognitive constructs that are subject to epistemic uncertainty, i.e., uncertainty that is an artifact of a lack of knowledge, partial informedness, and ignorance. Opinions about the solutions or constraints of a problem emerge earlier than hard evidence can be produced. Thus, relying on opinions to provide advice to reinforcement learning agents results in more agile and efficient reinforcement learning methods. For this, epistemic uncertainty has to be approached in a formal way in order to avoid unsound advice that would adversely affect the agent. Unfortunately, despite the fact that the importance of dealing with uncertainty in machine learning has been well-recognized [26], there is a lack of reinforcement learning methods that treat epistemic uncertainty as a first-class citizen in guided reinforcement learning. Specifically, methods for managing epistemic uncertainty are missing [26].
Traditional probability cannot handle constructs such as ignorance or partial knowledge, and can lead to unsound formalization of the advice. For example, when the advisor lacks knowledge about a particular situation, the only way to model the advice is to assign a 0.5 probability to it. A probability of 0.5 means that given a statement x captured in the advice (e.g., “there is high reward at location [2, 3]” in a grid world), x and not (x) are equally likely. This clearly does not represent ignorance and, thus, could mislead the RL agent. In general, forcing users to express their opinions by traditional probability could lead to unreliable conclusions. The RL agent needs to know that the advice is based on lacking knowledge in order to consider it with proper weight. It would be, thus, preferable to be able to say “I don’t know exactly” or “I’m not sure” within a piece of advice, and couple these modifiers with a particular statement (which might exhibit traits of traditional probability).
Subjective logic [28] is an extension of probabilistic logic [1], in which users can express subjective opinions by quantified parameters of belief and certainty. Opinions are formed from a belief component and an uncertainty component. In statistics and economics, the uncertainty component of subjective logic is often called second-order probability [22], and is represented in terms of a probability density function over first-order probabilities. Unfortunately, the complex framework of subjective logic limits its applicability. Most techniques that rely on it resort to simplifications, e.g., assuming that humans can express their opinions in terms of mathematical abstractions [10]; or emulating human opinion by objectively measured metrics, taking away the core subjective element of the approach [52].
Goal and research questions. Our goal is to evaluate the feasibility and utility of guiding RL agents by opinions—i.e., through advice that is subject to belief uncertainty—captured in constructs of subjective logic. To meet our goal, we formulate the following research questions.
  1. RQ1.
    How can one use opinions to guide reinforcement learning agents?
    By answering this research question, we aim to identify methods to formally externalize uncertain opinions and introduce them into the agent’s policy as guiding pieces of information.
  2. RQ2.
    How does opinion-based guidance affect the performance of reinforcement learning agents?
    By answering this research question, we aim to identify relevant changes in key performance metrics of opinion-guided RL agents.
We develop a method in which opinion is formulated through the sound mathematical foundations of subjective logic, and fused into the policy of reinforcement learning agents. Subsequently, we evaluate the performance of the advised agent through a series of experiments. Our results indicate that opinions, even if uncertain, improve the performance of reinforcement learning agents, resulting in higher rewards, more efficient exploration, and a better reinforced policy. Since opinions emerge earlier and more easily than hard evidence could be produced, opinion-guided reinforcement learning offers a more efficient and economical alternative to traditional guided reinforcement learning approaches.
Contributions. The main contribution of this work is a modeling approach for guiding reinforcement learning through subjective opinions, i.e., advice that is subject to belief uncertainty. We investigate various flavors of reinforcement learning algorithms and show where and how to augment them with hints expressed via subjective logic. Since the human-in-the-loop can quickly become a bottleneck in human-machine collaboration due to its expensive reward function [13], we also recommend architectural patterns to implement opinion-guided RL. Through detailed experiments, we show that the performance of RL agents improves with added human input and improves again with added uncertainty and disbelief information. Our approach fosters more efficient collaboration between human and machine agents [37] while improving the performance of RL agents.
Structure. The rest of this paper is structured as follows. In Sec. 2, we present the running example we use throughout the paper to illustrate key concepts, and to drive the evaluation of our approach. In Sec. 3, we briefly review the background topics relevant to our work and review the related work. In Sec. 4, we outline our approach to modeling human belief and certainty in reinforcement learning. In Sec. 5, we evaluate our approach under various advice types, sources, and learning parameters. In Sec. 6, we discuss the results and their implications and outline a set of open challenges. In Sec. 7, we draw the conclusions and outline future work.
2 Running example
To illustrate the principles of our approach, we rely on the following running example of a self-driving car through a frozen lake. The example is analogous to Open AI’s Gym’s Frozen Lake environment. [^1] https://gymnasium.farama.org/environments/toy_text/frozen_lake/
Figure 1: The Frozen Lake running example
Fig. 1 visualizes the Frozen Lake; an ordinary grid world, on which an agent (a self-driving car) moves one square at a time. The agent begins its exploration at the top leftmost tile (Start). Its aim is to reach the bottom rightmost tile (Goal). The agent traverses the Frozen Lake one tile at a time by choosing to move up, down, left, or right. There are three types of tiles the agent may encounter: several frozen tiles and holes, and one goal tile. Encountering these tiles has different consequences, as shown in Tab. 1.
Table 1: Reward structure
Event Consequence Reward
Stepping on a frozen tile Nothing happens; the agent is allowed to continue the traversal of the environment. 0
Stepping in a hole The task ends unsuccessfully. -1
Encountering the goal The task is completed successfully. 1
The goal of the agent is to learn the optimal path from the top leftmost tile to the bottom rightmost tile under the given reward structure. Through trial-and-error, the agent gradually learns which actions are beneficial in a specific situation. For example, in cell (2, 2), the agent should not move to the right, because this action would result in stepping into a hole and terminating the exploration unsuccessfully, with a reward of -1. Instead, in cell (2, 2), the agent should move down, because cell (3, 2) is a frozen tile, and it it allows the agent to move closer to the goal, which is in cell (3, 3).
Eventually, the agent will develop a policy, which will allow it to find the goal faster and in a more reliable fashion (i.e., not falling into holes).
3 Background and related work
Here, we provide a brief overview of the background of our work and review the related approaches.
3.1 Reinforcement learning
Figure 2: Reinforcement learning – conceptual overview [48]
The RL setting (Fig. 2) as described by Sutton and Barto [48], involves an agent that sequentially explores its environment, and uses feedback in the form of rewards to learn the optimal control of said environment. Typically, the RL process is formalized by finite horizon Markov decision processes [44], defined as a four-tuple $\langle\mathcal{S},\mathcal{A},P,r_{t}(s,a)\rangle$. Here, $\mathcal{S}$ denotes the set of states the agent may observe from the environment. $\mathcal{A}$ is the set of actions the agent may take in the environment. $\mathcal{R}\subset\mathbb{R}$ is the set of rewards, and $r_{t}(s,a)\in\mathcal{R}$ is the numeric reward received for choosing action $a\in\mathcal{A}$ while in state $s\in\mathcal{S}$ at time $t$. $P=p(s^{\prime},r|s,a)$ defines the dynamics of the environment, that is, the probability of the environment transitioning to state $s^{\prime}\in\mathcal{S}$ and producing reward $r\in\mathcal{R}$ given the agent is in state $s\in\mathcal{S}$ and performs action $a\in\mathcal{A}$. At each time step $t$, the agent observes a state $S_{t}\in\mathcal{S}$, and carries out an action $A_{t}\in\mathcal{A}$. At the next time step, $t+1$, the environment produces a new state $S_{t+1}$ as well as a reward $R_{t+1}\in\mathcal{R}$ based on the environment dynamics defined by $P$.
The agent uses a mapping from states to actions, called the policy $\pi(a|s)$, which gives the probability of the agent taking action $a\in\mathcal{A}$ given state $s\in\mathcal{S}$. The agent’s sequence of actions, as well as the states and rewards produced by the environment define the trajectory $\tau=(S_{0},A_{0},R_{1},...,S_{T-1},A_{T-1},R_{T})$ where $T$ is the time of termination. By sampling many trajectories through acting in the environment, the goal of the agent is to act according to an optimal policy $\pi_{*}$, that maximizes the sum of rewards, called the expected return, defined as $G_{t}=R_{t+1}+R_{t+2}+...+R_{T}$.
3.1.1 Various flavors of reinforcement learning
While all RL algorithms aim maximize expected return, various methods achieve this goal in different ways. Value-based methods aim to maximize either the state-value function $v_{\pi}(s)=\mathbb{E}_{\pi}[G_{t}|s]$, the action-value function $q_{\pi}(s,a)=\mathbb{E}_{\pi}[G_{t}|s,a]$, or both. Value-based approaches are deterministic because they select actions greedily when maximizing the value function; however, this might lead to under-exploration. In contrast, the goal of policy-based methods is to learn a parameterized policy that maximizes the reward function in every state, without the use of value functions. Policy-based methods explore the state space more thoroughly than value-based methods, but produce estimates with more noise, potentially resulting in unstable learning processes. Finally, actor-critic methods provide a trade-off between value-based and policy-based methods, in which the actor implements a policy-based strategy, and the critic criticizes the actions made by the actor based on a value function. Different methods offer various benefits and perform with unique utility in specific learning problems.
In particular, policy gradient is a subset of the aforementioned policy-based methods. In policy gradient, the policy is parameterized through a parameter vector $\boldsymbol{\theta}$. For reasonably sized action/state spaces, the policy may be parameterized as numerical preferences for all state-action pairs $h(s,a,\boldsymbol{\theta})\in\mathbb{R}$, and actions can then be chosen using a soft-max distribution. This is called discrete policy gradient, and often, a matrix-like representation of a look-up table is used. In contrast, the policy may be instead parameterized via a neural network such that $\boldsymbol{\theta}$ is a vector of connection weights of the network, or linear in features, which defines deep policy gradient. In either case, the policy is defined as $\pi(a|s,\boldsymbol{\theta})$, the probability of taking action $a$ in state $s$ with parameter $\boldsymbol{\theta}$. The policy may be parameterized in any manner, subject to the condition that it is differentiable with respect to $\boldsymbol{\theta}$. This way, $\boldsymbol{\theta}$ may be updated to maximize some scalar performance metric $J(\boldsymbol{\theta})$.
3.2 Subjective logic
To formalize human guidance in reinforcement learning, we rely on subjective logic. Subjective logic [28] is an extension of probabilistic logic [1], in which users can express subjective opinions by quantified parameters of belief and certainty. Quantifying subjective opinions is an improvement over traditional logic systems as it promotes opinions to first-class citizens in a systematic fashion. Well-informed opinions can be early indicators of emerging new knowledge and often, opinions are the only available insights in engineering. Subjective logic has seen growing applications, such as in model-driven software engineering, e.g., by Burgueño et al [10], where it is used to model belief uncertainty [5] in domain models. As well, it has been used in knowledge graphs as per Navarrete and Vallecillo [42] to allow reasoning about graph databases enriched with uncertainty.
3.2.1 Formal underpinnings
Given a boolean predicate $x$, a binomial opinion regarding the truth of $x$ is given by $\omega_{x}=(b_{x},d_{x},u_{x},a_{x})$, where $b_{x}$ represents the belief in the truthfulness of $x$; $d_{x}$ represents the disbelief in the truthfulness of $x$; $u_{x}$ quantifies the degree of epistemic uncertainty or uncommitted belief concerning $x$; and $a_{x}$ represents the base rate, i.e., the prior probability of $x$ in the absence of (dis)belief. For each parameter, $0\leq b_{x},d_{x},u_{x},a_{x}\leq 1$. These parameters satisfy the following conditions.
$b_{x}+d_{x}+u_{x}=1$ (1)
$P_{x}=b_{x}+a_{x}u_{x}.$ (2)
Equation 2 expresses the projected probability of an opinion, effectively transforming subjective opinions into the probability domain, where $P(x)$ means the probability that boolean predicate $x$ holds. It is clear that as uncertainty $u_{x}$ increases, projected probability $P_{x}$ is closer to base rate $a_{x}$. In contrast, as uncertainty $u_{x}$ decreases, projected probability $P_{x}$ is closer to that of the belief parameter $b_{x}$. Both Equations 1 and 2 are important invariants that we rely on throughout the paper, particularly in Sections 4 and 5.
Informed opinions work well at scale. That is, the more experts that express their opinions, the higher the credibility of a collective opinion. Collective opinions can be inferred by semantically sound fusion operators [30]. A fusion operator is a function $f:\Omega\times\Omega\to\Omega$ that maps a pair of subjective opinions onto a new, joint subjective opinion. Various fusion semantics exist, and the right operator must be chosen based on its fit for purpose. A detailed account of fusion operators is given in Sec. 4.3.3.
Example Formulating opinion about the Frozen Lake
Let $x$ be a boolean predicate regarding frozen lake as follows; $x$: action $a$ is beneficial when in state $s$. Now, consider moving right while in cell $[1,0]$ of Fig. 1. This would cause the agent to move to a hole tile, resulting in the task to ending unsuccessfully. In this case, a reasonable opinion would be that the action of moving right is likely not beneficial while in cell $[1,0]$, corresponding to $\omega=(0.0,1.0,0.0,0.25)$, meaning there is absolute disbelief that moving right while in cell $[1,0]$ is beneficial.
3.2.2 Challenges in employing subjective logic
The high expressive power of subjective logic comes with the added challenge of calibrating its parameters, particularly uncertainty $u$ and base rate $a$. There are two schools of thought to address this problem. In permissive techniques, setting uncertainty and belief-disbelief are deferred to the user [10, 27], and it is assumed that these parameters will become available at some point. However, it is usually not explained when, how, and who has to set these parameters, leading to a limited applicability of subjective logic. In restrictive techniques, users are not asked to parameterize the framework. Instead, parameters are controlled by external measures [29, 39]. Unfortunately, this takes away subjective elements from subjective logic. Margoni and Walkinshaw [39] and Walkinshaw and Shepperd [52] use statistical evidence to set values, employing subjective logic to analyze empirical studies and experimental results, respectively. In regards to base rate $a_{x}$, some sources opt to set this value via statistical evidence, such as in Jøsang and Bondi [29], Margoni and Walkinshaw [39]. Navarrete and Vallecillo [42] allow users to directly indicate the base rate, and Walkinshaw and Shepperd [52] assume $a_{x}=0.5$.
In our approach, we alleviate to cognitive load on the advisor by deriving base rate $a$ from the structure of the problem; and calibrate uncertainty $u$ by appropriate distance metrics between advisor and advised state (Sec. 4.2).
3.3 Related work
The closest work to ours is the one by Guo et al [24], who combine feedback from several human trainers to create a reliable reward for the RL agent. SL is used to model the agent’s opinion or trust of the human trainer. Converting this opinion into a probability yields a measure of the trainer’s trustworthiness, and after the agent utilizes human feedback, the agent’s opinion about the trainer’s is updated. In contrast, our work uses SL the other way around: to model the advisor’s opinion about the environment, and uses that opinion to drive policy shaping and improve the agent’s performance.
Other related works focus either on guiding reinforcement learning agents but without the use of subjective logic (Sec. 3.3.2) and employing subjective logic in reinforcement learning but not for guidance (Sec. 3.3.1).
3.3.1 Subjective logic in reinforcement learning (but not for guidance)
The combination of RL and SL is present in a number of publications. Notable work includes Zennaro and Jøsang [56], an initial inquiry into using SL in multi-armed bandits, a simplified RL problem. In this work, the agent forms an opinion over the available actions, and if the agent receives a positive reward for an action, the opinion about said action can be updated. Subjective logic to capture uncertainty in multi-armed bandit like problems is extended in Wang and Grace [53]. Similarly, Guo et al [25] uses a Deep Reinforcement Learning (DRL) multi-agent in smart farm monitoring to collect data from sensors on cattle- opinions about each animal’s condition is computed via this sensor data. In [58], RL is used to learn a low dimensional representation of safe regions of complex dynamical systems. Safety of states is estimated using DSAF (Discretized Safety Assessment Function), which is described with SL. In Guo et al [25], SL is used to model and compute uncertainty in a DRL framework used to identify the intents of tweets. In Zhao et al [57], opinions are used to represent edges in a graph network, and for unknown edges, a set of best paths is determined via DRL. This work proposes 3 different DRL models, each with reward given based on different types of uncertainty that are computed via SL. Similarly, Zhao et al [57] uses SL to represent users within an Online Social Network. Different types of users are initialized with particular opinions, which can be updated via SL fusion operators to enhance DRL based Competitive Influence Maximization. While these works combine RL and SL, our work differs by modeling both RL policy and user advice in SL, and using this model to guide the SL agent via human opinion.
3.3.2 Guidance in Reinforcement Learning (without SL)
While reinforcement learning algorithms are self-sufficient-that is, they can run without external intervention-there is a large body of work related to incorporating external human knowledge into these algorithms to improve performance. There are a vast number of differing approaches for for incorporating human advice into reinforcement learning without the use of subjective logic. Najar and Chetouani [41] survey the field of reinforcement learning with human advice, which they define as “teaching signals that can be communicated by the teacher to the learning system without executing the task”. They define a taxonomy of methods, first differentiating general advice from contextual advice. General advice is characterized by general constraints and construction, such as if-then rules  [38], or action plans  [51]. In contrast, contextual advice is subdivided into feedback, and guidance. Feedback methods aim to be evaluative and corrective, such as in Cruz et al [15], and Dai et al [17]. Reward shaping is a particular form of feedback, as seen in Tenorio-Gonzalez et al [49], where feedback is given directly and translated into rewards. Alternatively, guidance involves using human input to bias the exploration strategy, such as probabilistic, early, and importance advising as detailed in Cruz et al [16]. Policy shaping is a particular subset of guidance methods, where guidance is used to bias the exploration strategy of the agent. This has been used in several works, such as Griffith et al [23], Knox and Stone [32], Kessler Faulkner and Thomaz [31], Brawer et al [8], and Cederborg et al [11]. Following this taxonomy, our work falls under the category of guidance, and specifically, policy shaping, as we aim to bias the agent’s exploration by infusing the policy with human opinion by way of subjective logic.
Other methods not explicitly mentioned in the survey include demonstration or imitation based learning, such as in Celemin et al [12], Pertsch et al [43], and Wu et al [55]. Such training protocols have limitations, however, as the human must know how to properly demonstrate the task, and may become the bottleneck in learning leading to increased time to train the agent. Our work facilitates quicker human-machine interaction as the agent does not require continuous human interaction through the process in order to learn. Alternatively, Alshiekh et al [2] propose the use of shielding, in which a shield is placed pre-exploration, listing safe actions for the agent to take if it tries to traverse an unsafe region. The shield may instead be placed during exploration if the agent decides to take an unsafe action. Through the use of subjective logic, our work instead provides more semantically rich information to the agent rather than a list of alternative actions.
4 Approach to guiding reinforcement learning agents by opinions
In this section, we present our approach to opinion-guided reinforcement learning, with an emphasis on modeling belief uncertainty for guiding purposes. We argue that subjective logic is an appropriate formalism to encode opinions, and therefore, opinions should be captured in terms of subjective logic. However, as explained in Sec. 3, subjective logic is not intuitive to human advisors, and therefore, we recommend supporting the human advisor with a domain-specific language (DSL) [46] to express their opinions. Calibrating the uncertainty parameter u of subjective logic is particularly challenging.
Thus, as shown in Fig. 3, in our approach, we construct opinions in two steps and use them as the guiding advice in a subsequent step. First, (1) the advisor (or multiple advisors) provide(s) their advice through a suitable DSL, without the notion of uncertainty. Subsequently (2), the advice is translated to an opinion by factoring in the level of uncertainty u and base rate a of the opinion. Uncertainty is calibrated by a suitable external function, and the base rate is obtained from the structure of the reinforcement learning problem at hand. Opinions are formulated in terms of subjective logic (certainty domain). To shape the agent’s policy with the opinion, the policy is translated to subjective logic as well, and (3) fused with the opinion within the mathematically sound framework of subjective logic. Eventually (4), the fused opinion is translated back to the probability domain as the shaped policy.
Figure 3: Overview of the approach
We now elaborate on each of these steps in detail, using the running example of Sec. 2 to illustrate concepts. (Subsections 4.14.4 correspond to steps 14 in Fig. 3.)
4.1 Providing advice
In the first step of the approach (1), the advisor provides their advice about the benefits of the agent being in a specific state, as perceived by the advisor. Advice is a subjective construct of the advisor and expresses beliefs about a statement, e.g., the reward that might be received for occupying a particular state. Advice can be expressed at any time, given sufficient support for interaction. In this work, we assume that advice is provided before the exploration begins; however, our approach can be extended to interactive reinforcement learning scenarios too [12].
In our approach an advice $\alpha$ is given by a mapping $\alpha:s\in S^{\prime}\mapsto v$, where $S^{\prime}\subseteq S$ is the known subset of the total problem space, $s\in S^{\prime}$ is a specific state within the known problem space, and $v$ is a value that expresses the benefit of the particular state in a suitable language. Since it is not always feasible to survey the whole problem space, we only assume that the advisor provides a set of advice w.r.t. $S^{\prime}\subseteq S$. For example, in Sec. 2, a partially informed advisor could formulate the advice “Cell (1,1) [a hole] is likely not beneficial”. Here, $s$ refers to cell [1,1] in the grid world, and “not beneficial” is the value of occupying cell [1,1]. This advice is then used to update the state-action pairs that lead to the advised cell from neighboring cells, thereby shaping the policy.
Unfortunately, the notion of value in the advice is not quantitative enough as of yet. Thus, advice needs to be formulated in a suitable domain-specific language (DSL) [46] that either allows for expressing quantitative advice intuitively or quantifies the advice in the background. DSLs are languages tailored to a specific problem and offer a restricted set of syntactic elements to express statements about the problem domain. This allows for working with domain concepts, e.g., expressing the value of occupying a cell in the frozen lake, rather than having to express the reward function in the usual terms of reinforcement learning. Working with domain concepts, in turn, narrows the gap between reinforcement learning and human cognition, substantially improving the usability of the approach.
Example The GridWorld DSL for the Frozen Lake problem
For the purposes of this article, we define the DSL in Listing 1 to provide advice about the grid world of the running example (Fig. 1). An example set of advices expressed in the DSL about the running example (Fig. 1) is shown in Listing 2.
1<AdviceList> ::= <Advice>+
2<Advice> ::= <AdviceLocation> ‘,‘ <AdviceValue>
3<AdviceLocation> ::= ‘[‘[0-9]+, [0-9]+‘]‘
4<AdviceValue> ::= ^-?[0-2]+
Listing 1: GridWorld grammar in EBNF notation
1[1,1], -2
2[1,3], -2
3[0,3], -1
4[3,3], +2
Listing 2: Example set of opinions for the problem in Fig. 1
The value of the advice represents the advisor’s idea about how beneficial the given cell can be for the agent. In our example DSL, the value is an integer that ranges between -2 and 2. In the example in Listing 2, the respective list of advice corresponds to the following cells of the running example (Fig. 1): two holes on the grid, which are not beneficial ([1,1], -2; and [1,3], -2); the goal, which is beneficial ([3,3], +2); and the frozen tile that is not as hazardous as a hole but still rather disadvantageous ([0,3], -1) due to the number of holes in its vicinity.
It is important to note that the above DSL is just an example, and alternative DSLs for different problems could be developed. However, our approach is general enough to accommodate any DSL.
Next, we show how advice is translated into the certainty domain as an opinion.
4.2 Translating advice to opinion
To treat advice by formal means, we map them onto opinions of subjective logic. As shown in Fig. 3 (2), this requires calibrating the level of uncertainty u at which the advice was communicated and setting base rate a (Sec. 4.2.1); and subsequently, compiling the advice at the specific level of uncertainty into an opinion by calculating the remaining parameters of subjective logic: disbelief d and belief b (Sec. 4.2.2).
4.2.1 Calibrating the base rate (a) and the level of uncertainty (u)
Calibrating base rate
Base rate a is the prior probability in the absence of belief or disbelief. In the context of reinforcement learning, the base rate represents the default likelihood of the agent taking a specific action in a specific state. Thus, given a set of actions A, the base rate is given as
$a=\frac{1}{|A|}.$ (3)
In the grid world of the running example, the base rate is $0.25$ in each state because the agent can choose from four actions to move to the next state. (If the agent is situated in an edge or corner cell and chooses an action that would lead out of the grid world, the agent stays in place.)
A similar approach is suggested by Zennaro and Jøsang [56, Sec 3.1].
Calibrating uncertainty
As explained in Sec. 3, there are two classes of methods to calibrate uncertainty u: manual (deferred to the advisor) and automated (calibrated by an appropriate metric). Since human advisors might find it challenging to quantify the uncertainty of their advice, in our approach, we recommend automated derivation of the uncertainty metric.
Uncertainty can be derived, for example, from an appropriate distance metric, assuming that uncertainty increases with distance. Distance can be a pure topological concept (e.g., a Manhattan distance in a grid world) or a more abstract concept, such as trace distances between design models [21], or a quantification of domain-expertise. Formally, distance is a function between an advisor $h$, and the location of the advice $a$ is given as $d:h\times a\mapsto\mathbb{R}^{+}$.
In the running example, we use a two-dimensional Manhattan distance (i.e., $n=2$) to estimate uncertainty:
$d(p,q)=\sum_{i=1}^{n}|p_{i}-q_{i}|.$
By this metric, the uncertainty of each opinion can be calibrated automatically, alleviating the cognitive load on the advisor.
4.2.2 Compiling advice into opinion by computing disbelief (d) and belief (b)
Once uncertainty $u$ has been calibrated, disbelief $d_{i}$ and belief $b_{i}$ of opinion $\omega_{i}$, for advice $\alpha_{i}$ can be computed. Given $0\leq u\leq 1$, the remainder of opinion weights has to be distributed between $b_{i}$ and $d_{i}$ in accordance with Equation 1. The calculation method of belief $b_{i}$ and disbelief $d_{i}$ for the $j$th item in the $n$-point scale used for expressing advice values, in ascending order of confidence from least to most confident, are given as follows.
$b_{i}$ $=\frac{j-1}{n-1}\times(1-u)~{}|~{}j\in\{1..n\}$ (4)
$d_{i}$ $=(1-u)-b_{i}.$ (5)
The corresponding algorithm is shown in Algorithm 1. Here, $\frac{j-1}{n-1}$ is the weight between b and d to split the remaining opinion weight of $1-u$.
Eventually, using Equations 35, the opinion is given as
$\omega_{i}=\left(\frac{j-1}{n-1}\times(1-u),\frac{n-j}{n-1}\times(1-u),u,\frac{1}{|A|}\right).$
(Here, we also used that (5$\leftarrow$(4) reduces to $\frac{n-j}{n-1}\times(1-u)$.)
Algorithm 1 Compiling advice into opinion
$Advice~{}\alpha$
$a$triangleright inferred base rate
$u$triangleright automatically calibrated uncertainty
$n$triangleright advice scale length
$Opinion~{}\omega$
$\omega.a\leftarrow a$
$\omega.u\leftarrow u$
$\omega.b\leftarrow(order(\alpha.value)-1)/(n-1)\times(1-\omega.u)$
$\omega.d\leftarrow 1-(\omega.b+\omega.u)$
return $\omega$
Three example mappings are shown in Tab. 2. For example, advice value $+1$ is the element number $4$ in the advice scale. At uncertainty level $u=0.2$, the remaining opinion weight is $1-0.2=0.8$; which is distributed between b and d in a $\frac{j-1}{n-1}=\frac{3}{4}$ weight towards b. That is, $b_{4}=0.8\times\frac{3}{4}=0.6$; and consequently, $b_{4}=0.2$.
Table 2: Example mappings of confidence levels onto the belief dimension of SL at different degrees of uncertainty $u$
Advice value Order j u=0.0 u=0.2 u=0.5 u=0.833
b d b d b d b d
-2 1 0.00 1.00 0.0 0.8 0.000 0.500 0.000 0.167
-1 2 0.25 0.75 0.2 0.6 0.125 0.375 0.042 0.125
0 3 0.50 0.50 0.4 0.4 0.250 0.250 0.084 0.084
+1 4 0.75 0.25 0.6 0.2 0.375 0.125 0.125 0.043
+2 5 1.00 0.00 0.8 0.0 0.500 0.000 0.167 0.000
Example From advice to opinions in the Frozen Lake problem
Given an advisor situated in the bottom-left corner of the field (Fig. 4), we model the uncertainty of a piece of advice by the distance between the advisor and the cell the opinion pertains to. For simplicity, we use a normalized Manhattan distance and consider the bottom-left cell at $u_{(3,0)}=0.0$ and the top-right corner at $u_{(0,3)}=1.0$ uncertainty; and consider a linear relationship between uncertainty and distance. Given the Manhattan distance of 6 between the two corners, each unit of distance from the advisor increases the uncertainty by $1/6=0.166$.
Given the map of uncertainties in Fig. 4, the list of advice in Listing 2 translates to subjective opinions as shown in Tab. 3.
Figure 4: Visual intuition of the advisor’s limited knowledge, subject to epistemic uncertainty. Uncertainty grows with the topological distance. As uncertainty increases above $u=0.5$ (dashed arc), the advice is closer to a coin toss rather than truly informed knowledge. Using the distance between the Advisor and the cell as the main influencer of uncertainty ($\Delta\sim u$), the uncertainty of specific cells can be calibrated.
Table 3: The four advices and their subjective opinion equivalents of the running example
Cell Advice SL
u b d a
[1,1] -2 0.500 0.000 0.500 0.250
[1,3] -2 0.833 0.000 0.167 0.250
[0,3] -1 1.000 0.000 0.000 0.250
[3,3] +2 0.500 0.500 0.000 0.250
4.3 Policy shaping by opinion fusion
After having formed subjective opinions from advice, the advisor’s input is ready to be used to shape the agent’s policy. As shown in Fig. 3 (3), this requires accessing the agent’s policy (Sec. 4.3.1), transforming it from the probability domain to the certainty domain (Sec. 4.3.2), and within the framework of subjective logic, fusing it with the advisor’s opinion (Sec. 4.3.3).
4.3.1 Policy
The product of the agent’s exploration process in reinforcement learning is the policy $\pi$ that maps states to actions. Formally, policy $\pi$ is given by a mapping $\pi=S\times A\rightarrow P$, where S is the set of states accessible to the agent; A is the set of actions the agent can take, and P is a probability in the usual sense, i.e., $P=[0,1]$. Policy $\pi(a|s)$ defines the conditional probability of choosing action $a\in A$ in state $s\in S$. It follows that $\forall s\in S,a\in A:0\leq\pi(a|s)\leq 1$ always holds. Furthermore, the default policy is initialized as $\forall s\in S,a_{i},a_{j}\in A,i\neq j:\pi(a_{i}|s)=\pi(a_{j}|s)$. That is, before exploration, the probability of choosing any action in a given state is equal.
Example Default policy in the Frozen Lake problem
Tab. 4 shows the default policy of the running example. As shown, in each state (here: each cell), there are four actions available for the agent (left, down, right, up). In the initial state of the agent, each action is equally likely to be taken in each state. This probability is 0.25 due to the number of possible actions. Through trial and error, the agent will modify these probabilities.
Table 4: Default policy of the Frozen Lake running example
Left Down Right Up
[0,0] 0.25 0.25 0.25 0.25
[0,1] 0.25 0.25 0.25 0.25
[3,2] 0.25 0.25 0.25 0.25
[3,3] 0.25 0.25 0.25 0.25
4.3.2 Transforming the policy into the certainty domain
To be able to shape the policy, we transform it from the probability domain to the certainty domain. That is, the probabilities of the policy (e.g., in Tab. 4) are translated to subjective opinions through the following mapping:
$f_{P\rightarrow\Omega}(\pi):S\times A\rightarrow\Omega~{}~{}|~{}~{}P.$
That is, mapping $f(\pi)$ assigns a subjective opinion $\omega\in\Omega$ of choosing action $a\in A$ in state $s\in S$, given the probability $p\in P$ of choosing action $a\in A$ in state $s\in S$ under policy $\pi$. As explained in Sec. 3.2, probability $p$ is translated to an opinion $\omega$ as follows:
$\omega:p\mapsto(p,1-p,0,p).$ (6)
This definition also shows that base rate $a$ (i.e., the last element of the $\omega$ tuple) is equal to the default probability of choosing an action, aligned with Equation 3.
The corresponding algorithm is shown in Algorithm 2. For clarity, we refer to a policy expressed in the probability domain as $\pi_{p}$, and to a policy expressed in the certainty domain as $\pi_{c}$.
Algorithm 2 Transforming a policy in the probability domain ($\pi_{p}$) to a policy in the certainty domain ($\pi_{c}$)
$\pi_{p}$triangleright Policy in the probability domain
$\pi_{c}$triangleright Policy in the certainty domain
for Each state $s\in S$ do
     for Each action $a\in A$ do
         $p\leftarrow\pi_{p}[s,a]$
         $\pi_{c}[s,a]\leftarrow(p,1-p,0,p)$
     end for
end for
return $\pi_{c}$
Example Default policy of the Frozen Lake problem in the certainty domain
Tab. 5 shows the default policy in the certainty domain transformed from the default policy in the probability domain shown in Tab. 4.
As shown, the previous probabilities of $p=0.25$ are now expressed as opinions, following Equation 6 as $\omega=(0.25,1-0.25,0,0.25)=(0.25,0.75,0,0.25)$.
Table 5: Default policy of the Frozen Lake running example in the certainty domain
Left Down Right Up
[0,0] (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25)
[0,1] (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25)
[3,2] (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25)
[3,3] (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25)
4.3.3 Policy shaping by opinion fusion
After the policy has been transformed to an opinion-based representation, the advisors’ opinions can be combined with it. Combined (joined) opinions are obtained through fusing single opinions. By fusion, we mean a mapping $\Omega\times\Omega\rightarrow\Omega$ that produces a subjective opinion from other subjective opinions.
Locating the opinions to be fused
In our approach, the advisor provides advice about the value of occupying a specific state. This information affects the agent’s decision in states from which the advised state is reachable. To unambiguously formalize this step, we need some definitions.
Definition 1 (Neighboring states)
States $s_{i},s_{j}\in S$ are said to be neighbors if there exists a state-action pair in the policy for which $\pi(a|s_{i})\mapsto s_{j}$, i.e., choosing action $a\in A$ in state $s_{i}$ will transition the agent to state $s_{j}$.
The neighborhood $N$ of a state is the set of all its neighbors.
Definition 2 (Neighborhood (of a state))
$N(s_{i}\in S)=\bigcup_{s_{j}\in S}\exists a\in A:\pi(a|s_{i})\mapsto s_{j}$.
In the running example, the state space is topological; thus, trivially, neighboring states are the neighboring cells of a given cell.
Theorem 1
Introducing advice $\alpha$ about state $s_{i}$ to policy $\pi$ is achieved by $\forall a\in N(s_{i}).A:\omega(\alpha)\odot\omega(a)$.
That is, the opinion formed from the advice ($\omega(\alpha)$) is fused with every opinion formed from actions ($\omega(a)$) that lead from its neighboring states ($N(s_{i})$) to $s_{i}$.
Choosing a fusion operator
The seminal work of Jøsang [28] defines an array of fusion operators and classifies them according to situational characteristics to facilitate the selection of the most appropriate operator. Here, we rely on the Belief Constraint Fusion (BCF) operator, which is an appropriate choice when agents and advisors have already made up their minds and will not seek compromise. Since the agent and advisors formulate opinions about the problem independently from each other, it is fair to assume that they will, indeed, not seek compromise.
Under Belief Constraint Fusion, a fused opinion ${\omega}^{\odot}=({b}^{\odot},{d}^{\odot},{u}^{\odot},{a}^{\odot})$ is calculated from the overlapping beliefs (called harmony) and non-overlapping beliefs (conflict) of individual opinions.
$\textit{Harmony}=b_{1}u_{2}+b_{2}u_{1}+b_{1}b_{2};$ (7)
$\textit{Conflict}=b_{1}d_{2}+b_{2}d_{1}.$ (8)
Finally, the parameters of joint opinion ${\omega}^{\odot}$ are calculated as follows.
${b}^{\odot}$ $=\frac{\textit{Harmony}}{(1-\textit{Conflict})};$ (9)
${d}^{\odot}$ $=1-({b}^{\odot}+{u}^{\odot});$ (10)
${u}^{\odot}$ $=\frac{u_{1}u_{2}}{(1-\textit{Conflict})};$ (11)
${a}^{\odot}$ $=\frac{a_{1}(1-u_{1})+a_{2}(1-u_{2})}{2-u_{1}-u_{2}}$ (12)
Of course, different fusion operators can be chosen as well. A detailed account of fusion operators is given by Jøsang [28].
Example Shaped policy of the Frozen Lake problem in the certainty domain
Tab. 6 shows the impact of a single advice [1,1$\rightarrow$-2 on the default policy shown in Tab. 5. As defined in Theorem 1, shaping the policy impacts the neighbors of the advised state.
Locating opinions to be fused. From Definition 2, $N[1,1]=\{[0,1],[1,0],[1,2],[2,1]\}$, with the respective actions of $\{Down,Right,Left,Up\}$ leading to $[1,1]$. The corresponding opinions will be fused with the opinion formed from the advice.
Fusing opinions. As shown in Tab. 3, advice [1,1$\rightarrow$-2 translates to $\omega(\alpha)=(0.00,0.50,0.50,0.25)$. In the default policy, each opinion is given by $\omega(a)=(0.25,0.75,0.00,0.25)$. The fused opinion under Belief Constraint Fusion is calculated as follows.
Harmony $=b_{1}u_{2}+b_{2}u_{1}+b_{1}b_{2}=0.00\times 0.00+0.25\times 0.50=0.125.$
Conflict $=b_{1}d_{2}+b_{2}d_{1}=0.00\times 0.75+0.25\times 0.50=0.125.$
${b}^{\odot}$ $=\frac{\textit{Harmony}}{(1-\textit{Conflict})}=\frac{0.125}{1-0.125}=0.143.$
${u}^{\odot}$ $=\frac{u_{1}u_{2}}{(1-\textit{Conflict})}=\frac{0.50\times 0.00}{1-0.125}=0.$
${d}^{\odot}$ $=1-({b}^{\odot}+{u}^{\odot})=1-(0.143+0)=0.857.$
${a}^{\odot}$ $=\frac{a_{1}(1-u_{1})+a_{2}(1-u_{2})}{2-u_{1}-u_{2}}=\frac{0.25\times(1-0.50)+0.25\times(1-0.00)}{2-0.50-0.00}=0.25.$
Thus, the resulting fused opinion ${\omega}^{\odot}=(0.143,0.857,0.000,0.250)$, as highlighted in Tab. 6.
Table 6: Fused policy of the Frozen Lake running example after advice [1,1$\rightarrow$-2
Left Down Right Up
[0,0] (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25)
[0,1] (0.25, 0.75, 0.0, 0.25) (0.143, 0.857, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25)
[1,0] (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.143, 0.857, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25)
[1,2] (0.143, 0.857, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25)
[2,1] (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.143, 0.857, 0.0, 0.25)
[3,3] (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25) (0.25, 0.75, 0.0, 0.25)
4.4 Transforming fused policies from the certainty domain to the probability domain
As the final step of the approach (Fig. 34), the fused policy is transformed back from the certainty domain to the probability domain to be used by the agent. This is achieved in two steps: translating opinions to probabilities and, subsequently, normalizing probabilities in each state.
Transformation to the probability domain
The transformation to the probability domain is the inverse transformation of what has been explained in Sec. 4.3.2. As explained in Sec. 3.2 and defined by Jøsang [28], opinion $\omega$ is translated to a probability $p$ as follows:
$p:\omega\mapsto b+au.$ (13)
Normalization
Since the actions available for the agent in a given state form a complete probability space, the following invariant must always hold in a valid policy $\pi$.
$\forall s\in S:\sum_{a\in A}p(a|s)=1.$ (14)
Thus, we apply the usual normalization to scale the sum of probabilities to 1 as follows.
$\forall s\in S,a\in A:p(a|s):=p(a|s)\times\frac{1}{\sum_{a\in A}p(a|s)}.$ (15)
Example Shaped policy of the Frozen Lake problem in the probability domain
Calculating probabilities. We take the fused policy in Tab. 6 and for each tuple, we apply Equation 13. Tab. 7 shows the result of this transformation and the new policy that considers advice [1,1$\rightarrow$-2. Consider, e.g., the Down action in state [0, 1]. Here, $\omega=(0.143,0.857,0.0,0.25))$, and from Equation 13, $p=b+au=0.143+0.25\times 0.0=0.143$ follows.
As seen by the highlighted values, the advice affects the actions of neighboring states that lead to state [1,1], i.e., [0,1] - Down, [1,0] - Right, [1,2] - Left, [2,1] - Up. Tab. 7(a) and Tab. 7(b) show the policy before and after normalization, respectively.
Normalization. The resulting probabilities in row [0,1] violate invariant Equation 14 as $\sum_{a\in A}p(a|s=[0,1])=0.893$. Thus, we normalize by Equation 15 and multiply each probability by $\frac{1}{\sum_{a\in A}p(a|s=[0,1])}$, resulting in the probability vector (0.28, 0.16, 0.28, 0.28), as shown in the [0,1] row of Tab. 7(b).
Full example. To round out the running example, we show the full policy after considering every piece of advice in Listing 2. Tab. 8 shows the the new probabilities before (Tab. 8(a)) and after (Tab. 8(b)) normalization. As seen, some advice does not change the probabilities of the policy much, e.g., in the [0,3] - Down row of Tab. 8(a) (0.217, previously 0.250); while other pieces of advice have substantial impact, e.g., in [3,2] - Right (0348, previously 0.250). This is due to the uncertainty of the advice we modeled by the distance between the advisor and the advised state. In the running example, the advisor was situated in the bottom-left corner, and therefore, as a consequence, their advice about the distant [1,3] cell impacted the Down action of cell [0,3] less than the advice about the closer [3,3] cell that impacted the Right action of cell [3,2].
Table 7: Fused policy of the Frozen Lake running example in the probability domain after advice [1,1$\rightarrow$-2 affecting the actions of neighboring states [0,1], [1,0], [1,2], [2,1] that lead to state [1,1]
(a) Before normalization
Left Down Right Up
[0,0] 0.25 0.25 0.25 0.25
[0,1] 0.25 0.143 0.25 0.25
[1,0] 0.25 0.25 0.143 0.25
[1,2] 0.143 0.25 0.25 0.25
[2,1] 0.25 0.25 0.25 0.143
[3,3] 0.25 0.25 0.25 0.25
(b) After normalization
Left Down Right Up
[0,0] 0.25 0.25 0.25 0.25
[0,1] 0.28$\uparrow$ 0.16$\downarrow$ 0.28$\uparrow$ 0.28$\uparrow$
[1,0] 0.28$\uparrow$ 0.28$\uparrow$ 0.16$\downarrow$ 0.28$\uparrow$
[1,2] 0.16$\downarrow$ 0.28$\uparrow$ 0.28$\uparrow$ 0.28$\uparrow$
[2,1] 0.28$\uparrow$ 0.28$\uparrow$ 0.28$\uparrow$ 0.16$\downarrow$
[3,3] 0.25 0.25 0.25 0.25
Table 8: Complete fused policy of the Frozen Lake running example
(a) Before normalization
Left Down Right Up
[0,0] 0.25 0.25 0.25 0.25
[0,1] 0.25 0.143 0.25 0.25
[0,2] 0.25 0.25 0.250 0.25
[0,3] 0.25 0.217 0.25 0.25
[1,0] 0.25 0.25 0.143 0.25
[1,2] 0.143 0.25 0.217 0.25
[1,3] 0.25 0.250 0.25 0.25
[2,1] 0.25 0.25 0.25 0.143
[2,3] 0.25 0.400 0.25 0.217
[3,2] 0.25 0.25 0.400 0.25
[3,3] 0.25 0.25 0.25 0.25
(b) After normalization
Left Down Right Up
[0,0] 0.25 0.25 0.25 0.25
[0,1] 0.28$\uparrow$ 0.16$\downarrow$ 0.28$\uparrow$ 0.28$\uparrow$
[0,2] 0.25 0.25 0.25 0.25
[0,3] 0.259$\uparrow$ 0.223$\downarrow$ 0.259$\uparrow$ 0.259$\uparrow$
[1,0] 0.28$\uparrow$ 0.28$\uparrow$ 0.16$\downarrow$ 0.28$\uparrow$
[1,2] 0.166$\downarrow$ 0.29$\uparrow$ 0.252$\uparrow$ 0.29$\uparrow$
[1,3] 0.25 0.25 0.25 0.25
[2,1] 0.28$\uparrow$ 0.28$\uparrow$ 0.28$\uparrow$ 0.16$\downarrow$
[2,3] 0.224$\downarrow$ 0.358$\uparrow$ 0.224$\downarrow$ 0.194$\downarrow$
[3,2] 0.217$\downarrow$ 0.217$\downarrow$ 0.348$\uparrow$ 0.217$\downarrow$
[3,3] 0.25 0.25 0.25 0.25
Figure 5: The result of policy shaping in the running example with the major changes highlighted. Red: decreased probability; green: increased probability.
Visualization of results. The final policy in Tab. 8(b) is visualized in Fig. 5, with the major changes highlighted. As seen, the negative advice about the hole in [1,1] resulted in a strong decrease in the corresponding probabilities of actions from neighboring cells, while the negative advice about the hole in [1,3] resulted in slightly smaller decrease due to higher uncertainty of the advice. Conversely, the positive advice about the goal in [3,3] coupled with high certainty resulted in substantial positive bias that will guide the agent around the hole to the goal. Note that this policy is only the input to the agent’s learning phase, and the preferences in each state have yet to be reinforced. For example, in state [0,2], there is no change based on the advice, but surely, the agent will learn not to move down toward a hole but rather, to the right.
Conclusion of RQ1
In Sec. 1, we posed our first research question: How can one use opinions to guide reinforcement learning agents? In this section, we presented a method for guiding RL agents through opinions.
[…]
frametitle= RQ1 , innertopmargin=2pt, frametitleaboveskip=-frametitlealignment= {mdframed} Our method shows that through a combination of subjective logic, appropriate domain-specific languages, and a suitable distance function, opinions can be introduced into reinforcement learning as the means of guidance. The sound mathematical framework of subjective logic allows for multiple sources of opinions to be uniformly fuse with the agent’s original policy, effectively achieving policy shaping through opinions. Decoupling the calibration of uncertainty allows for more streamlined and intuitive advising by human experts.
5 Evaluation
In this section, we empirically evaluate our approach and investigate the effects of opinion-based human advice on the performance of reinforcement learning agents.
5.1 Study design
We conduct the study in a series of experiments on a sufficiently scaled-up version of the Frozen Lake running example (Sec. 2). To find the most appropriate configuration of the running example that is challenging enough but still feasible for the unadvised agent, we first manually experiment with our framework (Sec. 5.1.1), and based on our experiences we finalize the scenario (Sec. 5.1.2) and the setup (Sec. 5.1.3).
5.1.1 Manual experimentation
Before we run extensive, long-running experiments, we first manually investigate the effect of map configuration on the minimum number of episodes required for the vanilla RL agent to learn appropriately. In general, the larger the map, the harder it is to explore it, and therefore, more episodes are required for the agent to start functioning appropriately. The placement of termination points (holes) is another factor that might increase the minimum number of required episodes.
To experiment with various map configurations in a reproducible way, we developed a map generator that is able to generate maps of arbitrary size and hole density and disperse holes on the map randomly, controlled by a user-defined seed.
Eventually, we choose a $12\times 12$ map with 20% hole density (which is the default density in OpenAI Gymnasium) under seed 63. The final map is shown in Appendix A. On this map, we observe that the behavior of the unadvised RL agent is appropriate after 5 00050005\,0005 000 episodes. Thus, we set 5 00050005\,0005 000 as the minimum number of episodes we use in our evaluation.
5.1.2 Scenario
The scenario is a scaled-up version of the running example. We use the frozen lake environment, as described in detail in Sec. 2. The size of the lake in our experiments is 12$\times$12 with 20% hole ratio (which is the default value in Gym’s Frozen Lake environment [^2] https://gymnasium.farama.org/environments/toy_text/frozen_lake/#arguments). We also ensure that the goal is reachable from the start in the map.
Map generation
To generate such a large map, we developed a map generator in Python that allows for generating maps of arbitrary size and hole ratio. Holes are distributed randomly on the lake. For replicability, the generator can be parameterized with a seed. The generator outputs an Excel file for human inspection and the standard graphical rendering of the generated environment. The Excel file is then used as input for each experiment execution.
5.1.3 Experimental setup
Each experiment is executed using the same setup as discussed below.
Parameters
The list of parameters is shown in Tab. 9.
Table 9: Parameters of the experiments
Parameter Value
Fixed parameters
RL method Discrete policy gradient
Number of episodes 10 000
Advice strategy Once, at the beginning of the experiment
Fusion operator BCF
Grid size 12
Variables
Agent type {Random, Non-adviced, Advised}
Source of advice {Synthetic (oracle), Human}
Synthetic advice type {Every known cell (“ALL”), Just the holes and goal (“HOLES&GOAL”)}
Human advice quota {5%, 10%}
Degree of uncertainty ($u$) {$0.2k~{}|~{}k\in{0..5}$}
Fixed parameters. In this work, we choose discrete policy gradient as the reinforcement learning method. Policy-based methods represent the policy explicitly, which allows for directly investigating the effects of advice on the policy.
Each experiment runs for a number of episodes, defined by the number of episodes parameter. An episode starts with a reset environment and the agent in the initial position. Within the same experiment, the policy is updated and not reset. An episode lasts until the agent either finds the goal or steps into a hole. The higher the maximum number of episodes, the more opportunities agents have to find the goal and reinforce a beneficial policy. After we observed in our manual experiments that the behavior of the unadvised RL agents is appropriate after 5 000 episodes, we decided to run our experiments for a comfortable 10 000 episodes.
The advice strategy stipulates advice to be provided once and that it is done at the beginning of the experiment. Alternatively, more interactive advice strategies could be explored, but this is outside the scope of the current study.
The fusion operator is also fixed. We used the Belief Constraint Fusion (BCF) operator for simplicity. BCF is an appropriate choice when agents have already made up their minds and will not seek compromise [28]. Since the source of the advice formulates their opinions about the map independently from the RL agent, it is fair to assume that compromise will not be sought.
Variables. We execute experiments for each agent type to compare their performance. The random agent randomly samples from the available actions. Non-advised agents maintain a policy. Advised agents maintain a policy, and in the first step, they fuse the advice into their initial policy. In each case, the initial policy is the one in which the probability of choosing an action is uniformly distributed.
We execute experiments with different sources of advice. In the synthetic mode, a hypothetical and idealized oracle formulates its advice. We use two advice types in the synthetic mode: advice about every cell (“ALL”); and advice about the holes and the goal (“HOLES&GOAL”). In the human advice mode, a human formulates realistic advice based on partial knowledge. We tie the number of advices to the size of the problem and use 5% and 10% of the number of cells as the advice quota.
Both in synthetic and human advice mode, we execute experiments with different degree of uncertainty ($u$) that sweeps through the [0.0, 1.0] interval in 0.2 increments. Here, $u=0$ reduces to classic probability, and $u=1$ represents a uniform distribution [28].
Performance metric and analysis methods
Following community best practices, we use the cumulative reward as the metric of performance, i.e., the total reward $R_{T}=\sum_{i=0}^{T}r_{t+1}$ collected by the agent throughout an episode. We observe the amount of cumulative reward and its dynamics, i.e., how early and rapidly the cumulative reward increases under different settings.
Execution
We execute experiments using each combination of parameters defined above. We run each experiment 30 times to achieve sufficient statistical power. The results of experiments are saved as csv files.
We run the experiments on standard office equipment with the following parameters: Apple Mac mini, equipped with a 3.6 GHz Quad-Core Intel Core i3 CPU, running Mac OS Sonoma 14.14.1 with 8 GB 2667 MHz DDR4 of memory, using Python 3.11.5. This setup is satisfactory for our purposes as we are not interested in runtime performance.
5.1.4 Threats to validity and limitations
Construct validity
Observations of improved exploration performance may be artifacts of fortunately aligned problems (here: maps) or a few fortunate pieces of advice rather than artifacts of systematic effects. To mitigate this threat to validity, first, we experimented with different map layouts and observed the same performance trends. To facilitate consistent manual experimentation, we developed a map generator that generates map layouts in a reproducible way. Second, we ran our scaled-up experiments with a variety of settings, including both human and synthetic advisors.
Internal validity
Evaluating the approach with human advisors inevitably poses threats to internal validity. Specifically, maturation and testing might have influenced our results as the authors experimented with advising strategies and advice values. For example, cells that are not holes or goals are labeled freely with -1, 0, or 1, and a human might learn how to misuse these values to achieve a specific behavior by the agent. We mitigated these threats by defining clear advise score rules, which are also used by the synthetic advisor.
External validity
We chose only one flavor of reinforcement learning—a policy-based method—to demonstrate our approach, and we did not provide evidence that our approach can be applied to other reinforcement learning methods. However, we are reasonably confident that our methods translate to reinforcement learning methods where the policy is not explicitly represented. For example, in value-based methods, advice can be fused into the expected cumulative rewards in a given state, similar to the approach presented in this article. Nonetheless, in future work, we will validate this assumption.
Limitations
We used a relatively simple advising strategy, in which advice is provided before the exploration. It is reasonable to assume that more interactive advising modes might shed light on additional benefits and challenges in guided reinforcement learning. However, we are reasonably confident that the scope of this article was sufficient to draw key takeaways that hold in more interactive reinforcement learning setups as well. Nonetheless, we will experiment with interactive advising modes in future work.
5.2 Results
In the following, we report the results of our experiments. First, we have a look at the performance of synthetic advisor (Sec. 5.2.1), and subsequently, we review the performance of real human advice (Sec. 5.2.2). In both cases, we first investigate the performance in terms of the cumulative reward, and then we look at the final policy.
As the key takeaway of our work, Tab. 10 highlights that even uncertain advice at $u=0.6$ improves the performance of the agent. We observe mixed results with very uncertain advice at $u=0.8$; here, the volume of advice by the synthetic agent somewhat compensates for the low certainty, and keeps the performance slightly above that of the unadvised agent, but a human advisor with limited advice quota and such high uncertainty fails to improve over the unadvised agent. Thus, $u=0.6$ seems to be the level of uncertainty at which opinion-guided agents consistently outperform the unadvised agent.
Table 10: Cumulative rewards by experiment. (Bold is best at the given level of certainty. Red is below the performance of the unadvised agent.)
Synthetic Human
All Holes 10% 5%
Advice@u=0.0 8 778.033 9 045.933 8 541.633 6 674.333
Advice@u=0.2 7 584.600 6 904.033 6 226.367 3 527.967
Advice@u=0.4 4 702.267 3 023.000 4 345.033 1 539.600
Advice@u=0.6 3 102.833 1 789.500 2 373.800 855.933
Advice@u=0.8 1 638.000 909.967 285.400 527.867
Advice@u=1.0 489.600 183.833 351.833 461.633
Random 0.100
No advice 823.367
5.2.1 Performance of guidance with synthetic advisor
Cumulative reward
Fig. 6 shows the cumulative rewards of the agent with a synthetic advisor under two different advice strategies. Each plot shows the random agent and the agent without advice, compared to advised agents at five levels of uncertainty ($u=0.0,0.2,0.4,0.6,0.8,1.0$).
Linear scale
(a) Advice strategy: ALL
(b) Advice strategy: HOLES&GOAL
Log scale
(c) Advice strategy: ALL
(d) Advice strategy: HOLES&GOAL
Figure 6: Cumulative reward with a synthetic advisor
Clear performance improvement in advised agents. As the most important takeaways, we observe that (i) advised agents clearly outperform the unadvised agent and that (ii) performance improves with the level of certainty of the advice. This can be seen from the lines corresponding to lower uncertainty consistently being over lines corresponding to higher uncertainty. As expected, the random agent hardly accumulates any reward as it follows a random trajectory without learning.
Improved performance manifests in two forms. First, the cumulative reward is higher and increases at a consistently higher pace in advised and certain agents, as indicated by the advised lines over others in the linear and log scales, respectively. Second, cumulative reward increases earlier and reaches a gradual slope sooner in advised and certain agents than in the unadvised agent, indicated by the earlier rise of advised lines in the log scale charts.
Even substantially uncertain advice makes a difference. Particularly demonstrated in the $u=0.6$ case, agents with low-certainty advice can still outperform the unadvised agent. In very uncertain cases, i.e., $u=0.8$ and $u=1.0$, the performance of the advised agent is similar to that of the unadvised agent. The extensive “all” advise strategy compensates for low certainty and renders the advised agent in the $u=0.8$ case more performant than the unadvised agent.
Learning dynamics: more rapid learning. The log scale charts further show the two main phases of the learning process. In the first phase, approximately between episodes 0–2 000, we see a steep increase, indicating fast learning. In the second phase, approximately after episode 2 000, the slopes become more gradual, indicating a slowed-down collection of reward and a likely convergence to the maximum potential performance. Performance differences remain consistent throughout the experiment. The effect of advice, thus, is shifting the exponentially growing initial phase earlier in the learning process.
Advice strategies: difference shows only in low-certainty cases. There is not much difference between the two advice strategies in high-certainty cases, as shown by the very similar green ($u=0.0$) and red ($u=0.2$) lines in Fig. 5(a) and Fig. 5(b). In more uncertain cases ($u>0.2$), there is a noticeable performance deterioration in the holes-and-goal strategy, as neither line reaches the same eventual cumulative reward as in the “all” strategy.
Effects on the policy
(a) No advice
Advice strategy: ALL
(b) $u=0.4$
(c) $u=0.0$
Advice strategy: HOLES&GOAL
(d) $u=0.4$
(e) $u=0.0$
Figure 7: Heatmap visualization of the final learned policy with a synthetic advisor
Fig. 7 visualizes the policy as a heatmap. The cells of each heatmap correspond to the cells of the Frozen Lake problem; arrows represent the action with the highest probability; and shading corresponds to the probability. Empty cells show unexplored states.
Better reinforced policy. One of the two key improvements is indicated by the darker shades of cells compared to the unadvised agent. Darker shades correspond to higher probabilities of the indicated action in a cell and an overall better reinforced beneficial policy. By providing advice, the agent navigated through the state space more efficiently and had time (episodes) to explore crucial cells, e.g., the ones around holes and the goal. The “holes&goal” strategy reinforces some key cells better than the extensive “all” strategy, e.g., around the goal and around the starting point.
More thorough exploration. The other key improvement is shown by the higher number of cells explored in advised strategies as both the “all” and “holes&goal” advice strategies cover more cells compared to the unadvised agent. This is possible due to the advice driving the agent in crucial situations, resulting in fewer failed attempts. The “holes&goal” strategy allows to explore more states at higher levels of certainty than the extensive “all” strategy (cf. Fig. 7(e) and Fig. 7(c)); however, this relation turns around in the lower-certainty case (cf. Fig. 7(b) and Fig. 7(d)).
Effects clear even at relatively low certainty. At $u=0.4$, we still see demonstrated improvements compared to the unadvised agent. That is, even fairly uncertain advice has a clear positive impact on the policy.
5.2.2 Performance of guidance with human advisor
Cumulative reward
Fig. 8 shows the cumulative rewards of the agent with a human advisor under two different advice quotas.
Key performance improvement trends are retained. We see similar trends to those in Sec. 5.2.1. That is, advised agents outperform the unadvised agent even with a severely reduced advice quota. As a reminder, the “all” strategy provided advice about every cell; the “holes&goal” strategy provided advice about the holes and the goal, i.e., about 20% of the cells—here, the human operating under 10% and 5% quotas is still able to improve the agent’s performance. With 10% quota, even at $u=0.4$, the performance reaches that of the synthetic advisor.
The role of certainty increases as the advice quota is reduced. This trend is indicated by the distances between the lines both in the linear and log scales between 10% and 5%. At 10% advice quota, the performance differences between different levels of certainty are similar; at 5% advice quota, performance drops more prominently as certainty decreases.
Advice quotas: differences more pronounced in low-certainty cases. The difference between the two quotas is more obvious as certainty decreases. In the 5% case, the fully certain advice $u=0.0$ only reaches the level of somewhat certain advice $u=0.2$ of the 10% case. Above uncertainty level $u=0.2$, the 5% quota deteriorates substantially, marking the likely limit at which opinion-based guidance can be efficient.
Linear scale
(f) Advice quota: 10%
(g) Advice quota: 5%
Log scale
(h) Advice quota: 10%
(i) Advice quota: 5%
Figure 8: Cumulative reward with a human advisor
Effects on the policy
Fig. 9 visualizes the policy as a heatmap. The cells of each heatmap correspond to the cells of the Frozen Lake problem; arrows represent the action with the highest probability; and shading corresponds to the probability. Empty cells show unexplored states.
Key improvements are retained. We see similar trends to those in Sec. 5.2.1. The policy in advised agents is better reinforced, as indicated by the darker shades of blue, especially around holes and the goal. We see more explored states by advised agents, though the difference between moderately certain advice with low advice quota, and no advice is less substantial.
Certainty becomes more of a factor. In line with the observations about the cumulative reward in the human-advised mode, certainty becomes an important factor in advice. The policy under 5% advice quota and at $u=0.4$ shows minor improvements over the unadvised mode, while fully certain advice at $u=1.0$ still shows substantial differences.
(a) No advice
Advice quota: 10%
(b) $u=0.4$
(c) $u=0.0$
Advice quota: 5%
(d) $u=0.4$
(e) $u=0.0$
Figure 9: Heatmap visualization of the final learned policy with a human advisor
Conclusion of RQ2
In Sec. 1, we asked How does opinion-based guidance affect the performance of reinforcement learning agents? as our second research question. In this section, we designed and executed a study to investigate this research question.
[…]
6 Discussion
In this section, we discuss the results and some of the important implications.
6.1 The impact of advice – Discussion of the results
Even uncertain advice is useful
The key takeaway of this work is that even uncertain opinions improve the performance of the learning agent. Even with a human advisor, opinions as uncertain as $u=0.6$ guided agents outperform vanilla agents. Considering that opinions are formed much faster than hard evidence can be produced, opinion-based guidance can be an important improvement over traditional guidance mechanisms if opinions can be efficiently elicited from advisors.
To provide an intuition of what uncertainty at the level of $u=0.6$ means, e.g., in terms of the Frozen Lake problem, consider the two edge cases when an advisor is very sure about a state being (1) disadvantageous and (2) advantageous. Given that $u=0.6$, the remaining opinion weight is $0.4$, which has to be distributed over disbelief d and belief b. The first case would translate to an opinion $\omega_{1}=(0.4,0.0,0.6,0.25)$, and the second to $\omega_{2}=(0.0,0.4,0.6,0.25)$. Using the formulas from Sec. 3.2, the two opinions translate to $P_{1}=0.0+0.15=0.15$ and $P_{2}=0.4+0.15=0.55$, respectively. That is, the advisor has between 0.15–0.55 probability—good for a meager 40% wiggle room—to express the advantage of choosing a particular state. The advisor cannot express anything below 15% and above 55%, and yet, these values are still vastly superior to not giving advice without misleading the agent.
Finding trade-offs between certainty, exhaustiveness, and cost of advice
A key trend we observe is that performance deteriorates as uncertainty increases. In some cases, low-certainty advice might have a negative impact on the performance. Shown in Tab. 10, agents advised at $u=1.0$ perform below the vanilla agent, although Figures 5(c)-5(d) and 7(h)-7(i) show a narrowing gap between low-certainty advice and no-advice agents, indicating that low-certainty advice would likely catch up with no-advice agents. Thus, the safe level of uncertainty in our experiments is around $u=0.6$, and any advice beyond that level of uncertainty is not much of a help. Adopters of the approach should assess the limits of their own settings and decide about the accepted levels of uncertainty accordingly.
An interesting trend to observe with the synthetic advisor is that the “holes&goal” strategy reinforces some key cells better than the “all” strategy, e.g., around the goal and around the starting point. At the same time, the “holes&goal” strategy allows to explore more states at higher levels of certainty than the “all” strategy (cf. Fig. 7(e) and Fig. 7(c)); however, this relation turns around in the more uncertain cases (cf. Fig. 7(b) and Fig. 7(d)). It is plausible to conclude that selective and targeted advice strategies might lead to more pronounced differences in the shaped policy, and by allowing for an improved exploration rate, they transitively allow for a better-reinforced policy. Thus, adopters of our approach should look for trade-offs between exhaustiveness and cost of advice as the returns of facilitating advising the agents might diminish.
Ultimately, adopters should seek trade-offs between each of these quality properties: certainty, exhaustiveness of advice, and cost of advice. We foresee the identification and management of these trade-offs to become the key to the successful and practical adoption of our approach.
Synthetic vs. Human advice
Comparing synthetic and human advisors’ performance does not show substantial differences (Tab. 10). In particular, the human advisor with sufficient quota achieves performance similar to that of the synthetic advisor. In our experiments, the human with 10% advice quota (i.e., the number of pieces of advice is 10% of the number of all cells) achieved similar scores at $u=0.0$$u=0.6$ as the synthetic advisors that had 100% (“all”) and 20% (“holes&goal”) quota. This suggests that human creativity is superior to synthetic strategies, both to exhaustive style and strategically targeted ones. Thus, ours is another one in the long line of works that highlight the role of human cognition in computer-automated methods, such as machine learning. Since the human advisor can quickly become the bottleneck in reinforcement learning, adopters of our approach should consider implementing efficient advice strategies, particularly asynchronous ones in which human and machine agents can work independently.
6.2 Open challenges and research opportunities
Our approach opens up new challenges and research opportunities for prospective researchers.
Opinion-based guiding in value-based and deep reinforcement learning
In our approach, we opted for a policy-based RL algorithm because it allows for better investigation of the impact of opinion on the explicitly represented policy. We are fairly confident that our method translates to value-based techniques as well, but validation is left for future work. A fairly more complicated challenge is introducing guidance into deep reinforcement learning, where the policy is encoded in a neural network, and the lack of explainability might hinder the development of a sound guiding framework. With the demonstrated utility of deep reinforcement learning in numerous problems, such as the management of engineering systems [3] and simulator inference [19], its support by opinion-based guidance seems like a natural next step.
Interaction patterns for optimal guidance
We relied on one-time advice provided before exploration. However, more sophisticated advice dynamics have been used in guided reinforcement learning [41, 45]. We recommend researching more interactive patterns of advising. A known drawback of placing humans into the learning loop is the high cost of human reward signals, which are caused by the human stakeholders quickly becoming performance bottlenecks [33]. We see opportunities in opinion-based guidance in this aspect as opinions might be easier to formulate than hard evidence, alleviating the load on the human. Another interesting research opportunity our approach motivates is the investigation of push and pull mechanisms in advising reinforcement learning agents, in which the human can push advice toward the agent, and the agent can actively query the human for advice at any time. Such mechanisms would allow for the human to provide advice under limited quotas when advice matters the most.
Languages for expressing advice
A key advantage of opinion-based guidance is the rapid pace at which opinions emerge. To leverage this advantage to its fullest, intuitive languages are needed to express opinions. Domain-specific languages (DLSs) [46] have been widely used to capture expert opinions at high levels of abstraction, narrowing the gap between the modeling language and human cognition. However, DSLs typically operate with reduced syntax to remain truly specific to the problem domain, and therefore, automated methods for DSL engineering are in high demand [9]. This is true for guided reinforcement learning as well. We envision language workbenches that allow for the automated construction of DSLs through which advice can be provided by domain concepts. For example, in the Frozen Lake example, one might want to use descriptive quantifiers, such as “hazardous” or “beneficial” instead of relying on an integer scale. Unfortunately, there is little research on domain-specific languages for reinforcement learning, and existing languages mostly focus on programming reinforcement learning procedures [40] rather than expressing opinions at high cognitive levels. We see plenty of room for research targeting domain-specific modeling languages for reinforcement learning, inspired by the advancements in the model-driven engineering community [18, 35, 54].
7 Conclusion
In this article, we presented a method to guiding reinforcement learning agents by opinions—cognitive constructs subject to uncertainty. We have devised a formal, mathematically sound method for using opinions for policy shaping, based on subjective logic, and demonstrated that opinions, even at moderate-to-substantial levels of uncertainty help achieve better performance in reinforcement learning. Specifically, we observe improved cumulative rewards and faster convergence to the theoretical maximum performance in opinion-guided reinforcement learning agents.
Our method is particularly useful in situations where hard evidence is impractical or infeasible to produce, e.g., in problems with a scarce reward structure, where human creativity plays an important role. Such situations are, for example, problems with infeasible, inaccessible, hazardous, or costly states, which, consequently, cannot be efficiently explored by the reinforcement learning agent. In addition, problems, where the expertise of multiple distinct domain experts is required, will benefit from our approach as well, thanks to the sound fusion semantics provided by subjective logic.
Although demonstrated in a topological problem, our approach translates naturally to problems where the notion of distance is more abstract, e.g., in guided design-space exploration. Such directions are left for future work. Additional future work will focus on extending the method to different flavors of reinforcement learning, supporting more interactive advice modes, and evaluating our method in real industry settings.
Appendix A Study setup
Figure 10: The map used in the experiments. (12x12 seed 63)
1[11,11], +2
2[0,4], -2
3[0,5], -2
4[1,4], -2
5[1,5], -2
6[1,7], -2
7[1,8], -2
8[1,9], -2
9[1,11], -2
10[2,7], -2
11[3,0], -2
12[3,7], -2
13[4,4], -2
14[4,9], -2
15[5,5], -2
16[6,3], -2
17[6,4], -2
18[6,5], -2
19[6,7], -2
20[7,1], -2
21[7,3], -2
22[8,3], -2
23[9,1], -2
24[9,4], -2
25[9,6], -2
26[10,1], -2
27[10,4], -2
28[11,10], -2
Listing 3: Input advice with an opinion about every cell of Fig. 10. Corresponds to the HOLES&GOAL strategy in the experiments.
1[11,11], +2
2[0,4], -2
3[0,5], -2
4[1,4], -2
5[1,5], -2
6[1,7], -2
7[1,8], -2
8[1,9], -2
9[1,11], -2
10[2,7], -2
11[3,0], -2
12[3,7], -2
13[4,4], -2
14[4,9], -2
15[5,5], -2
16[6,3], -2
17[6,4], -2
18[6,5], -2
19[6,7], -2
20[7,1], -2
21[7,3], -2
22[8,3], -2
23[9,1], -2
24[9,4], -2
25[9,6], -2
26[10,1], -2
27[10,4], -2
28[11,10], -2
29[0,0], +1
30[0,1], +1
31[0,2], +1
32[0,3], +0
33[0,6], +0
34[0,7], +0
35[0,8], +0
36[0,9], +0
37[0,10], +1
38[0,11], +0
39[1,0], +1
40[1,1], +1
41[1,2], +1
42[1,3], +0
43[1,6], -1
44[1,10], -1
45[2,0], +0
46[2,1], +1
47[2,2], +1
48[2,3], +1
49[2,4], +0
50[2,5], +0
51[2,6], +0
52[2,8], -1
53[2,9], +0
54[2,10], +1
55[2,11], +0
56[3,1], +0
57[3,2], +1
58[3,3], +1
59[3,4], +0
60[3,5], +1
61[3,6], +0
62[3,8], +0
63[3,9], +0
64[3,10], +1
65[3,11], +1
66[4,0], +0
67[4,1], +1
68[4,2], +1
69[4,3], +0
70[4,5], -1
71[4,6], +1
72[4,7], +0
73[4,8], +0
74[4,10], +0
75[4,11], +1
76[5,0], +1
77[5,1], +1
78[5,2], +1
79[5,3], +0
80[5,4], -1
81[5,6], +0
82[5,7], +0
83[5,8], +1
84[5,9], +0
85[5,10], +1
86[5,11], +1
87[6,0], +1
88[6,1], +0
89[6,2], +0
90[6,6], -1
91[6,8], +0
92[6,9], +1
93[6,10], +1
94[6,11], +1
95[7,0], +0
96[7,2], -1
97[7,4], -1
98[7,5], +0
99[7,6], +1
100[7,7], +0
101[7,8], +1
102[7,9], +1
103[7,10], +1
104[7,11], +1
105[8,0], +1
106[8,1], -1
107[8,2], +0
108[8,4], -1
109[8,5], +1
110[8,6], +0
111[8,7], +1
112[8,8], +1
113[8,9], +1
114[8,10], +1
115[8,11], +1
116[9,0], +0
117[9,2], +0
118[9,3], -1
119[9,5], -1
120[9,7], +0
121[9,8], +1
122[9,9], +1
123[9,10], +1
124[9,11], +1
125[10,0], +0
126[10,2], +0
127[10,3], +0
128[10,5], +0
129[10,6], +0
130[10,7], +1
131[10,8], +1
132[10,9], +1
133[10,10], +0
134[10,11], +1
135[11,0], +1
136[11,1], +0
137[11,2], +1
138[11,3], +1
139[11,4], +0
140[11,5], +1
141[11,6], +1
142[11,7], +1
143[11,8], +1
144[11,9], +0
Listing 4: Input advice with an opinion about every cell of Fig. 10. Corresponds to the ALL strategy in the experiments.